# Trade-offs in Data Memorization via Strong Data Processing Inequalities

### v1

@@ 1.1 Our Contribution
We next consider a Boolean clustering problem. In this problem $\theta\in\{\pm 1\}^{d}$, negative examples are sampled from the uniform distribution over {±$\{\pm 1\}^{d}$, whereas positive examples are sampled from a product distribution over {±$\{\pm 1\}^{d}$ with mean $\lambda\cdot\theta$. Thus for $\theta$ chosen uniformly from {±$\{\pm 1\}^{d}$, a random positive example from $\mathcal{P}_{\theta}$ is coordinate-wise $\lambda$-correlated with $\theta$. The problem can be thought of as a Boolean analogue of the Gaussian setting above, and we obtain nearly tight (up to a log factor) upper and lower bounds for this problem. The bounds are similar to those in the Gaussian case for $n\leq\sqrt{d}$ but almost no memorization is needed when $n\geq\sqrt{d}\log d$. We note however that the analysis of the approximately dominating variable is more involved in this case and is derived from composition results in differential privacy.
Finally, we consider a sparse Boolean hypercube clustering problem introduced in (Brown et al., 2021). In this setting, the data distribution is defined by a pair $\theta=(S,Y)$, where $S\subseteq[d]$ is a subset of coordinates and $Y\in\{\pm 1\}^{S}$ are the values assigned to these coordinates. In the data distribution $\mathcal{P}_{\theta}$ negative examples are uniform over {±$\{\pm 1\}^{d}$, whereas the positive examples are uniform over the hypercube of all the points $x$ whose values in coordinates in $S$ are exactly $Y$: namely $x\in\{\pm 1\}^{d}$ that satisfy the conjunction $\bigwedge_{i\in S}(x_{i}=Y_{i})$. To sample $\theta$ we include each index in $S$ with probability $\approx 1/\sqrt{d}$ independently at random, and then assign to each coordinate a uniformly random value in $Y\in\{\pm 1\}^{S}$. As noted in (Brown et al., 2021), when $n=1$ this problem is identical to the Boolean clustering problem defined above. However, for larger $n$, we show it requires much less memorization. Specifically, we give a lower and upper bound (tight up to log factors) that are:
@@ 2 Formal Problem Setting
[…]
@@ 3 General Framework: SDPIs and Memorization
In this section, we will introduce the main machinery that allows us to derive excess memorization lower bounds via strong data processing inequalities (SDPIs). We start by recalling the definition of an SDPI.
@@
Suppose $(A,B)$ are $\sqrt{\rho}$-correlated uniform Boolean vectors, namely $A\sim\mathcal{U}(\{\pm 1\}^{d}),~{}B=\mathrm{Bin}_{\frac{1-\sqrt{\rho}}{2}}(A)$. Then $A,B$ satisfy the $\rho$-SDPI.
@@ 3.1 SDPIs imply Memorization
[…]
@@
[…]
@@ 4.1 Gaussian Clustering
  • Any algorithm $\mathcal{A}$ with constant $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{G}})\leq\alpha<\frac{1}{2}$ satisfies
    $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{G}})\geq\frac{1-\lambda^{2}}{\lambda^{4}n}\cdot(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)=\Omega\left(\frac{d}{n}\right)~{}.$
  • As long as $n\leq c\sqrt{d}$ for some sufficiently small absolute constant $c>0$, the lower bound above is nearly-tight: There is a learning algorithm $\mathcal{A}$ such that $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{G}})\leq 0.01$ and $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{G}})=O(d\cdot\log(\frac{d}{n})/n)$.
@@
To prove the second item, we rely on the ideas presented in Section 3. Particularly, we note that the information about $\theta$ contained in the dataset $X_{1},\dots,X_{n}\sim\mathcal{N}(\lambda\theta,(1-\lambda^{2})I_{d})$ is dominated by the empirical average $\hat{X}:=\frac{1}{n}\sum_{i\in[n]}X_{i}\sim\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$. Noting that $\lambda\theta\sim\mathcal{N}(0_{d},\lambda^{2}I_{d})$, we see that the variance of $\hat{X}$ in every direction is $\lambda^{2}+\frac{1-\lambda^{2}}{n}=\frac{1+\lambda^{2}(n-1)}{n}$. Therefore, by rescaling $\hat{X}$ by $\sqrt{n/(1+\lambda^{2}(n-1))}$, we get a Gaussian with unit variance $Z_{\theta}^{\mathrm{train}}=\sqrt{n/(1+\lambda^{2}(n-1))}\cdot\hat{X}$ which is a dominating variable for the dataset, whose coordinate-wise correlation with a fresh sample $X\sim\mathcal{N}(\lambda\theta,(1-\lambda^{2})I)$ is $\lambda^{2}\sqrt{n/(1+\lambda^{2}(n-1))}$. By Fact 3.2, this gives us a test/train SDPI with $\rho_{n}=\lambda^{4}n/(1+\lambda^{2}(n-1))$. The same dominating variable also proves a data generation $\tau_{n}$-SDPI by computing the coordinate-wise correlation between $\theta$ and $Z_{\theta}^{\mathrm{train}}$ as $\lambda\sqrt{n/(1+\lambda^{2}(n-1))}$, and therefore $\tau_{n}=\lambda^{2}n/(1+\lambda^{2}(n-1))$ once again by Fact 3.2. Plugging these SDPI coefficients into Theorem 3.7, we see that $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{G}})\gtrsim(1-\tau_{n})/\rho_{n}=\frac{1-\lambda^{2}n/(1+\lambda^{2}(n-1))}{\lambda^{4}n/(1+\lambda^{2}(n-1))}=\frac{1-\lambda^{2}}{\lambda^{4}n}\approx d/n$.
To prove the last item, we provide a simple low error algorithm that returns a low error classifier which is describable using $\widetilde{O}(d/n)$ bits, by using projections. To that end, considering once again the empirical average $\hat{X}\sim\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$, we project it onto $\mathbb{R}^{\ell}$ for $\ell\approx d/n$, obtain $\hat{X}^{[1:\ell]}$, and consider the linear classifier that projects onto $\mathbb{R}^{\ell}$ and takes the inner product with $\hat{X}^{[1:\ell]}$. The idea here is that Gaussian concentration arguments ensure that on one hand for $X\sim\mathcal{P}_{\theta}:~{}\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle\gtrsim\lambda^{2}\ell\approx C^{2}d^{1/2}/n$ with high probability, while on the other hand for $X\sim\mathcal{P}_{0}=\mathcal{N}(0_{d},I_{d}):~{}\langle X^{[1:\ell]},\hat{X}^{[1:\ell]}\rangle\lesssim\sqrt{\ell/n}\approx d^{1/2}/n$. Therefore the classifier achieves low error, and moreover, requires roughly $\ell$ bits to fully describe (ignoring logarithmic terms due to quantization). In particular, the classifier cannot contain more than $\widetilde{O}(\ell)$ bits of information about the dataset, proving the claimed memorization upper bound. ∎
@@ 4.2 Boolean Clustering
$\mathcal{P}_{\theta}~{}=~{}\mathrm{Bin}_{\frac{1-\lambda}{2}}(\theta)~{},~{}~{}~{}~{}~{}\theta\sim\Psi~{}=~{}\mathcal{U}(\{\pm 1\}^{d})~{}.$
@@
  • Any algorithm $\mathcal{A}:\mathcal{X}^{n}\to\mathcal{M}$ with $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq\alpha<\frac{1}{2}$ satisfies
    $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=\Omega\left(\frac{1-\lambda^{2}n\log\log|\mathcal{M}|^{d}}{\lambda^{4}n\log\log|\mathcal{M}|^{d}}(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)\right)~{}.$
    In particular, as long as $n\leq c\sqrt{d}$ for some sufficiently small absolute constant $c>0$, and $|\mathcal{M}|\leq\exp(d^{\tilde{C}})$ for some absolute constant $\tilde{C}>0$, then
    $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=\widetilde{\Omega}\left(\frac{1-\lambda^{2}n}{\lambda^{4}n}(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)\right)=\widetilde{\Omega}\left(\frac{d}{n}\right)~{}.$
  • The lower bound above, as well as the sample size condition, are both nearly-tight: On one hand, if $n\leq\sqrt{d}$ then there is a learning algorithm $\mathcal{A}$ such that $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq 0.01$ and $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=O(d/n)$. On the other hand, there is a learning algorithm $\mathcal{A}$ such that $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq 0.01$ and $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=O(d^{2}\exp(-n/\sqrt{d}))$, and so if $n\gtrsim\sqrt{d}\log d$ then $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq 1/\mathrm{poly}(d)$.
@@
The proof of Theorem 4.3 is similar in spirit to that of Theorem 4.1 as we previously sketched, yet the technical details in the proof of the memorization lower bound (second item) are more challenging. This follows from the fact that, as opposed to the Gaussian case, the boolean dataset $X_{1:n}\sim\mathrm{Bin}_{\frac{1-\lambda}{2}}(\theta)^{n}$ does not have a simple dominating variable. Instead, we use arguments related to advanced composition from the differential privacy literature, to argue that $X_{1:n}$ is statistically close to a post-processing of $Z_{\theta}^{\mathrm{train}}\sim\mathrm{Bin}_{\frac{1-\xi}{2}}(\theta)$ for $\xi\approx\sqrt{n}\lambda$, namely a single variable which is $\sqrt{n}$-times more correlated with $\theta$. We can then invoke Fact 3.3 and use our approximate reduction in Theorem 3.8 to obtain the required SDPIs, by noting that the coordinate-wise correlation of $\theta$ and $Z_{\theta}^{\mathrm{train}}$ is $\sqrt{\tau_{n}}=\xi$, whereas that of $Z_{\theta}^{\mathrm{train}}$ and $X\sim\mathrm{Bin}_{\frac{1-\lambda}{2}}(\theta)$ is $\sqrt{\rho_{n}}=\xi\lambda$. Consequently, up to negligible additive factors, we obtain $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\gtrsim\frac{1-\xi^{2}}{\lambda^{2}\xi^{2}}\approx\frac{1-\lambda^{2}n}{\lambda^{4}n}$, and under the assignment of $\lambda$ and assumption that $n\ll\sqrt{d}$, the latter simplifies to $d/n$.
The nearly-matching upper bounds are realized by an algorithm that computes the bit-wise majority vote over the sample. For the regime $n\lesssim\sqrt{d}$, the algorithm only computes the majority along the first $\ell\approx d/n$ coordinates, and returns a linear classifier in the projected space, similarly to the Gaussian case. Concentration arguments ensure that the algorithm has small error, while clearly requiring at most $\ell$ bits of memory, thus in particular no more than $\ell\approx d/n$ bits from the training set can be memorized. When $n\gtrsim\sqrt{d}\log(d)$, computing the majority vote in each coordinate reconstructs the parameter $\theta$ with very high confidence, so in this regime, accurate learning is possible with nearly zero excess memorization. The full proof appears in Appendix C.

### v2

@@ 1.1 Our Contribution
We next consider a Boolean clustering problem. In this problem $\theta\in\{\pm 1\}^{d}$, negative examples are sampled from the uniform distribution over $\{\pm 1\}^{d}$, whereas positive examples are sampled from a product distribution over $\{\pm 1\}^{d}$ with mean $\lambda\cdot\theta$. Thus for $\theta$ chosen uniformly from $\{\pm 1\}^{d}$, a random positive example from $\mathcal{P}_{\theta}$ is coordinate-wise $\lambda$-correlated with $\theta$. The problem can be thought of as a Boolean analogue of the Gaussian setting above, and we obtain nearly tight (up to a log factor) upper and lower bounds for this problem. The bounds are similar to those in the Gaussian case for $n\leq\sqrt{d}$ but almost no memorization is needed when $n\geq\sqrt{d}\log d$. We note however that the analysis of the approximately dominating variable is more involved in this case and is derived from composition results in differential privacy.
Finally, we consider a sparse Boolean hypercube clustering problem introduced in (Brown et al., 2021). In this setting, the data distribution is defined by a pair $\theta=(S,Y)$, where $S\subseteq[d]$ is a subset of coordinates and $Y\in\{\pm 1\}^{S}$ are the values assigned to these coordinates. In the data distribution $\mathcal{P}_{\theta}$ negative examples are uniform over $\{\pm 1\}^{d}$, whereas the positive examples are uniform over the hypercube of all the points $x$ whose values in coordinates in $S$ are exactly $Y$: namely $x\in\{\pm 1\}^{d}$ that satisfy the conjunction $\bigwedge_{i\in S}(x_{i}=Y_{i})$. To sample $\theta$ we include each index in $S$ with probability $\approx 1/\sqrt{d}$ independently at random, and then assign to each coordinate a uniformly random value in $Y\in\{\pm 1\}^{S}$. As noted in (Brown et al., 2021), when $n=1$ this problem is identical to the Boolean clustering problem defined above. However, for larger $n$, we show it requires much less memorization. Specifically, we give a lower and upper bound (tight up to log factors) that are:
@@ 1.3 Related Work
Lower bounds on memorization are also directly implied by lower bounds on differentially private algorithms. This follows from well-known results showing that the output of any $(\varepsilon,\delta)$ differentially private algorithm $\mathcal{A}$ has bounded mutual information with the dataset (McGregor et al., 2010; Bassily et al., 2018; Feldman and Steinke, 2018; Duchi and Rogers, 2019). Specifically, for any distribution $P$ over $\mathcal{X}$ and $X_{1:n}\sim P^{n}$, $I(\mathcal{A}(X_{1:n}),X_{1:n})=O(\varepsilon^{2}n+\delta n(H(P)+\log(1/\delta)))$ (see (Brown et al., 2021) for an outline of a special case of this statement). In particular, our lower bounds on memorization directly imply lower bounds on accurate differentially private algorithms for solving the learning problems we defined.
@@ 2 Formal Problem Setting
We abbreviate a sequence $X_{1},\dots,X_{n}$ by $X_{1:n}$. $I_{d}$ denotes the $d\times d$ identity matrix, $0_{d}$ denotes the $d$-dimensional zero vector, and we occasional omit the subscript when the dimension is clear from the context. Given a finite set $S$, we denote by $\Delta(S)$ the set of all distributions over $S$, and by $\mathcal{U}(S)$ the uniform distribution over $S$. We denote $X\perp\!\!\!\perp Y$ when $X,Y$ are independent random variables. $\|\cdot\|$ denotes the Euclidean norm, and $d_{\mathrm{TV}}(\cdot,\cdot)$ denotes the total variation distance (which when applied to random variables, is the distance between their corresponding distributions). For $x\in\{\pm 1\}^{d}$, $\mathrm{BSC}_{p}(x)$ denotes the product distribution over $\{\pm 1\}^{d}$ in which for $Y\sim\mathrm{BSC}_{p}(x)$, $\Pr[Y_{i}\neq x_{i}]=1-\Pr[Y_{i}=x_{i}]=p$ independently for every $i\in[d]$. The mapping from $x$ to $Y$ is usually referred to as the binary symmetric channel (BSC) in the context of data processing inequalities. Given random variables $X,Y,Z$, we denote by $H(X)$ the entropy of XX, [^2] We slightly abuse information theoretic notation by using it both for discrete and continuous random variables. In the latter case, definitions are with respect to the differential entropy. by I(X;Y):=H(X)H(X|Y)I(X;Y):=H(X)-H(X\,|\,Y) the mutual information between $X$ and $Y$, and by $I(X;Y\,|\,Z):=\mathop{\mathbb{E}}_{Z}[I(X\,|\,Z;Y\,|\,Z)]$ the conditional mutual information.
@@ 3 General Framework: SDPIs and Memorization
In this section, we will introduce the main machinery that allows us to derive excess memorization lower bounds via strong data processing inequalities (SDPIs). We start by recalling the definition of an SDPI. Specifically, we will consider so called input-dependent SDPIs in which the marginal distribution of the source random variable is fixed (Polyanskiy and Wu, 2024).
@@
Suppose $(A,B)$ are $\sqrt{\rho}$-correlated uniform Boolean vectors, namely $A\sim\mathcal{U}(\{\pm 1\}^{d}),~B=\mathrm{BSC}_{\frac{1-\sqrt{\rho}}{2}}(A)$. Then $A,B$ satisfy the $\rho$-SDPI.
@@ 3.1 SDPIs imply Memorization
$\mathrm{mem}_{n}(\mathbf{P},\alpha)\geq\frac{1-\tau_{n}}{\rho_{n}}C_{\alpha}-\mathrm{neg}_{n}~,\text{~~~where~~~}C_{\alpha}:=(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)~.$
@@
$=I(\mathcal{A}(X_{1:n});X_{1:n},\theta)-I(\mathcal{A}(X_{1:n});\theta)\qquad$ ${\color[rgb]{.25,.25,.25}\left[\text{chain rule}\right]}$
@@ 4.1 Gaussian Clustering
  • For $\alpha<1/2$,
    $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{G}},\alpha)\geq\frac{1-\lambda^{2}}{\lambda^{4}n}\cdot(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)~,$
    and, in particular, $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{G}})=\Omega\left(\frac{d}{n}\right)$.
  • The lower bound above is tight: There is a learning algorithm $\mathcal{A}$ such that $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{G}})\leq 0.01$ and $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{G}})=O(d/n)$.
@@
To prove the second item, we rely on the ideas presented in Section 3. Particularly, we note that the information about $\theta$ contained in the dataset $X_{1},\dots,X_{n}\sim\mathcal{N}(\lambda\theta,(1-\lambda^{2})I_{d})$ is dominated by the empirical average $\hat{X}:=\frac{1}{n}\sum_{i\in[n]}X_{i}\sim\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$. Noting that $\lambda\theta\sim\mathcal{N}(0_{d},\lambda^{2}I_{d})$, we see that the variance of $\hat{X}$ in every direction is $\lambda^{2}+\frac{1-\lambda^{2}}{n}=\frac{1+\lambda^{2}(n-1)}{n}$. Therefore, by rescaling $\hat{X}$ by $\sqrt{n/(1+\lambda^{2}(n-1))}$, we get a Gaussian with unit variance $Z_{\theta}^{\mathrm{train}}=\sqrt{n/(1+\lambda^{2}(n-1))}\cdot\hat{X}$ which is a dominating variable for the dataset, whose coordinate-wise correlation with a fresh sample $X\sim\mathcal{N}(\lambda\theta,(1-\lambda^{2})I)$ is $\lambda^{2}\sqrt{n/(1+\lambda^{2}(n-1))}$. By Fact 3.2, this gives us a test/train SDPI with $\rho_{n}=\lambda^{4}n/(1+\lambda^{2}(n-1))$. The same dominating variable also proves a data generation $\tau_{n}$-SDPI by computing the coordinate-wise correlation between $\theta$ and $Z_{\theta}^{\mathrm{train}}$ as $\lambda\sqrt{n/(1+\lambda^{2}(n-1))}$, and therefore $\tau_{n}=\lambda^{2}n/(1+\lambda^{2}(n-1))$ once again by Fact 3.2. Plugging these SDPI coefficients into Theorem 3.7, we see that $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{G}})\gtrsim(1-\tau_{n})/\rho_{n}=\frac{1-\lambda^{2}n/(1+\lambda^{2}(n-1))}{\lambda^{4}n/(1+\lambda^{2}(n-1))}=\frac{1-\lambda^{2}}{\lambda^{4}n}\approx d/n$.
To prove the last item, we consider an algorithm that computes $\tilde{\mu}=\hat{X}+\eta$ where $\hat{X}\sim\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$ is the empirical average and $\eta\sim N(0_{d},I_{d})$ is an independent isotropic Gaussian, and returns a linear classifier with respect to $\tilde{\mu}$. The idea is that Gaussian concentration arguments ensure that on one hand for $X\sim\mathcal{P}_{\theta}:\,\langle X,\tilde{\mu}\rangle\gtrsim\lambda^{2}\|\theta\|^{2}\gtrsim C^{2}\sqrt{d}$ with high probability, while on the other hand for $X\sim\mathcal{P}_{0}:\,\langle X,\tilde{\mu}\rangle\lesssim C\sqrt{d}$, and so setting a threshold of $C^{3/2}\sqrt{d}$ achieves low error. To bound the algorithm’s memorization, we analyze the conditional mutual information $I(\tilde{\mu},\hat{X}\mid\theta)=\mathop{\mathbb{E}}_{\theta\sim\Psi,u\sim\hat{X}}[D_{\mathrm{KL}}(u+\eta\,||\,\hat{X}+\eta)]$, and calculate that the latter is bounded by $O(d/n)$ using the formula for the KL divergence between Gaussians.
We remark that a previous version of this work provided a different algorithm, which exhibits a nearly-tight memorization upper bound (up to a log factor), yet it only applied when $n\lesssim\sqrt{d}$. This algorithm is based on a projection approach which we will also use for the Boolean application in the next section. For completeness, we include this algorithm as well as its analysis in the appendix as well. ∎
@@ 4.2 Boolean Clustering
$\mathcal{P}_{\theta}~=~\mathrm{BSC}_{\frac{1-\lambda}{2}}(\theta)~,~~~~~\theta\sim\Psi~=~\mathcal{U}(\{\pm 1\}^{d})~.$
@@
  • For $\alpha<\frac{1}{2}$,
    $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{B}},\alpha)=\Omega\left(\frac{1-8\lambda^{2}n\log(dn\log|\mathcal{M}|/C_{\alpha})}{\lambda^{4}n\log(dn\log|\mathcal{M}|/C_{\alpha})}C_{\alpha}\right)~,~~~C_{\alpha}:=(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)~.$
[…]
    $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{B}},\alpha)=\Omega\left(\frac{1-8\lambda^{2}n\log(dn/C_{\alpha})}{\lambda^{4}n\log(dn/C_{\alpha})}C_{\alpha}\right)~,$
    and $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{B}})=\widetilde{\Omega}\left(\frac{d}{n}\right)$.
  • The lower bound above is nearly-tight: On one hand, if $n\leq\sqrt{d}$ then there is a learning algorithm $\mathcal{A}$ such that $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq 0.01$ and $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=O(d/n)$. On the other hand, there is a learning algorithm $\mathcal{A}$ such that $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq 0.01$ and $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})=O(d^{2}n\exp(-n/\sqrt{d}))$, and so if $n\gtrsim\sqrt{d}\log(d)$ then $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{B}})\leq 1/\mathrm{poly}(d)$.
@@
The proof of Theorem 4.3 is similar in spirit to that of Theorem 4.1 as we previously sketched, yet the technical details in the proof of the memorization lower bound (second item) are more challenging. This follows from the fact that, as opposed to the Gaussian case, the boolean dataset $X_{1:n}\sim\mathrm{BSC}_{\frac{1-\lambda}{2}}(\theta)^{n}$ does not have a simple dominating variable. Instead, we use arguments related to advanced composition from the differential privacy literature, to argue that $X_{1:n}$ is statistically close to a post-processing of $Z_{\theta}^{\mathrm{train}}\sim\mathrm{BSC}_{\frac{1-\xi}{2}}(\theta)$ for $\xi\approx\sqrt{n}\lambda$, namely a single variable which is $\sqrt{n}$-times more correlated with $\theta$. We can then invoke Fact 3.3 and use our approximate reduction in Theorem 3.8 to obtain the required SDPIs, by noting that the coordinate-wise correlation of $\theta$ and $Z_{\theta}^{\mathrm{train}}$ is $\sqrt{\tau_{n}}=\xi$, whereas that of $Z_{\theta}^{\mathrm{train}}$ and $X\sim\mathrm{BSC}_{\frac{1-\lambda}{2}}(\theta)$ is $\sqrt{\rho_{n}}=\xi\lambda$. Consequently, up to logarithmic factors, we obtain $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{B}})\gtrsim\frac{1-\xi^{2}}{\lambda^{2}\xi^{2}}\approx\frac{1-\lambda^{2}n}{\lambda^{4}n}$, and under the assignment of $\lambda$ and assumption that $n\ll\sqrt{d}$, the latter simplifies to $\widetilde{\Omega}(d/n)$.
The nearly-matching upper bounds are realized by an algorithm that computes the bit-wise majority vote over the sample. For the regime $n\lesssim\sqrt{d}$, the algorithm only computes the majority along the first $\ell\approx d/n$ coordinates, and returns a linear classifier in the projected space. Concentration arguments ensure that the algorithm has small error, while clearly requiring at most $\ell$ bits of memory, thus in particular no more than $\ell\approx d/n$ bits from the training set can be memorized. When $n\gtrsim\sqrt{d}\log(d)$, computing the majority vote in each coordinate reconstructs the parameter $\theta$ with very high confidence, so in this regime, accurate learning is possible with nearly zero excess memorization. The full proof appears in Appendix C.